# On the Theory and Practice of GRPO: A Trajectory-Corrected Approach with Fast Convergence

### v1

@@ Abstract
Group Relative Policy Optimization (GRPO), recently proposed by DeepSeek, is a critic-free reinforcement learning algorithm for fine tuning large language models. It replaces the value function in Proximal Policy Optimization (PPO) with group normalized rewards, while retaining PPO style token level importance sampling based on an old policy. We show that GRPO update rule in fact estimates the policy gradient at the old policy rather than the current one. However, since the old policy is refreshed every few steps, the discrepancy between the two remains small limiting the impact of this bias in practice. We validate this through an ablation study in which importance sampling is entirely removed, and updates are instead performed using the gradient estimated at a fixed old policy across multiple optimization steps. Remarkably, this simplification results in performance comparable to standard GRPO.
@@ 2.1 Reinforcement Learning Setup
$J(\theta)=\mathbb{E}_{s_{T}\sim\pi_{\theta}}\left[r(s_{T})\right]=\sum_{s_{T}\in\mathcal{S}_{T}}\pi_{\theta}(s_{T}|s_{0})\,r(s_{T}),$
@@
$\theta_{n+1}=\theta_{n}+\eta_{n}\nabla_{\theta_{n}}\mathcal{J}(\theta_{n}),$
with learning rate sequence $\{\eta_{n}\}_{n\geq 1}$. Algorithms like PPO and GRPO build on this principle with various modifications to improve performance.
@@ 2.2 Review of GRPO
$\mathcal{L}_{\text{GRPO}}(\theta,\theta_{\text{old}},\theta_{\text{ref}})$ $=\frac{1}{|G|}\sum_{i=1}^{|G|}\sum_{t=1}^{T}\min\left\{w(s_{t}^{(i)},\theta,\theta_{\text{old}})A_{i},\overline{\operatorname{clip}}\bigl{(}w(s_{t}^{(i)},\theta,\theta_{\text{old}}),\epsilon_{\text{low}},\epsilon_{\text{high}}\bigr{)}A_{i}\right\}$
$\quad-\beta\,\mathrm{KL}\bigl{(}\pi_{\theta}\|\,\pi_{\theta_{\text{ref}}}\bigr{)},$ (4)
where the importance ratio $w(s_{t}^{(i)},\theta,\theta_{\text{old}})={\pi_{\theta}(s_{t}^{(i)}\mid s_{t-1}^{(i)})}\Big{/}{\pi_{\theta_{\text{old}}}(s_{t}^{(i)}\mid s_{t-1}^{(i)})},$ and the clipping function
$\overline{\operatorname{clip}}(x,\epsilon_{\text{low}},\epsilon_{\text{high}})=\begin{cases}1-\epsilon_{\text{low}},&x<1-\epsilon_{\text{low}},\\[4.0pt] x,&1-\epsilon_{\text{low}}\leq x\leq 1+\epsilon_{\text{high}},\\[4.0pt] 1+\epsilon_{\text{high}},&x>1+\epsilon_{\text{high}}.\end{cases}$
In original GRPO, the clipping thresholds used in the surrogate objective are symmetric, i.e., $\epsilon_{\text{low}}=\epsilon_{\text{high}}$. A subsequent study found that using asymmetric clipping ($\epsilon_{\text{low}}\neq\epsilon_{\text{high}}$) can lead to improved empirical performance, and accordingly renamed the modified algorithm as Decouple Clip and Dynamic Sampling Policy Optimization (DAPO).
In this paper, we adopt the same asymmetric setting as and base all our experiments on their DAPO configuration. However, for notational simplicity, we do not distinguish between GRPO and DAPO in the remainder of the paper.
@@
$\theta_{k+1}=\theta_{k}+\eta_{n}\widehat{\nabla}\mathcal{L}_{\text{GRPO}}(\theta_{k},\theta_{\text{old}},\theta_{\text{ref}}),$
where $\{\eta_{n}\}_{n\geq 1}$ is the learning rate  [^2] Here, $n$ denotes the outer loop index (i.e., the number of major update rounds). Within each inner loop—where the old policy $\pi_{\mathrm{old}}$ remains fixed—we assume the learning rate stays constant. See Algorithm 1 for details on the outer/inner loop structure. and $\widehat{\nabla}\mathcal{L}_{\text{GRPO}}(\theta,\theta_{\text{old}},\theta_{\text{ref}})$ denotes a mini-batch estimate of the gradient $\nabla\mathcal{L}_{\text{GRPO}}(\theta,\theta_{\text{old}},\theta_{\text{ref}})$ which can be written as
$\nabla\mathcal{L}_{\text{GRPO}}(\theta,\theta_{\text{old}},\theta_{\text{ref}})$ $=\underbrace{\frac{1}{|G|T}\sum_{i=1}^{|G|}\sum_{t=1}^{T}\nabla\left(\min\left\{w(s_{t}^{(i)},\theta,\theta_{\text{old}})A_{i},\overline{\operatorname{clip}}\bigl{(}w(s_{t}^{(i)},\theta,\theta_{\text{old}}),\epsilon_{\text{low}},\epsilon_{\text{high}}\bigr{)}A_{i}\right\}\right)}_{\emph{Gradient Term}}$
$\quad\underbrace{-\beta\nabla\left(\mathrm{KL}\bigl{(}\pi_{\theta}\|\,\pi_{\theta_{\text{ref}}}\bigr{)}\right)}_{\emph{Penalty Term}}.$ (5)
@@
Input: reward function $r(\cdot)$; initial parameters $\theta$; reference parameters $\theta_{\mathrm{ref}}$; group size $|G|;$ clipping threshold $\epsilon_{0}$; KL coefficient $\beta$;
inner iterations $K$; learning-rate schedule $\{\eta_{n}\}_{1:N}$
@@ 10 $A_{i}\leftarrow\dfrac{r_{i}-\mu_{G}}{\sigma_{G}+\delta}$;
14    $\mathcal{L}\leftarrow\frac{1}{|G|}\sum_{i,j}\left(\min\left\{w_{i,j}A_{i},\,\overline{\operatorname{clip}}\bigl{(}w_{i,j},\,\epsilon_{\text{low}},\epsilon_{\text{high}}\bigr{)}A_{i}\right\}\;+\beta\,D_{\mathrm{KL}}\!\Bigl{(}\pi_{\theta}\bigl{(}\cdot\mid s^{(i)}_{j-1}\bigr{)}\,\|\,\pi_{\theta_{\mathrm{ref}}}\bigl{(}\cdot\mid s^{(i)}_{j-1}\bigr{)}\Bigr{)}\right)$;
@@
16    $\theta\leftarrow\theta+\eta_{n}\,\hat{\nabla}\mathcal{L}$;
@@ 3 A Decomposition of GRPO’s Gradient Term
$\quad\text{\emph{Gradient Term}}\mathop{=}^{(*)}\frac{1}{|G|}\sum_{i=1}^{|G|}\sum_{t=1}^{T}\bm{1}_{\mathcal{B}(s_{t}^{(i)},\theta,\theta_{\text{old}})}w(s_{t}^{(i)},\theta,\theta_{\text{old}})\nabla\log\pi_{\theta}(s_{t}^{(i)}|s_{t-1}^{(i)})\cdot A_{i}$
${=}\frac{1}{\overline{\sigma}_{G}}\underbrace{\frac{1}{|G|}\sum_{i=1}^{|G|}\sum_{t=1}^{T}\bm{1}_{\mathcal{B}(s_{t}^{(i)},\theta,\theta_{\text{old}})}\nabla\log\pi_{\theta_{\text{old}}}(s_{t}^{(i)}|s_{t-1}^{(i)})\cdot\left(r_{i}-\overline{\mu}_{G}\right)}_{\widetilde{\nabla}J(\theta_{\text{old}})}$
$\quad+\underbrace{\frac{1}{\overline{\sigma}_{G}}\frac{1}{|G|}\sum_{i=1}^{|G|}\sum_{t=1}^{T}\bm{1}_{\mathcal{B}(s_{t}^{(i)},\theta,\theta_{\text{old}})}\frac{1}{\pi_{\theta_{\text{old}}(s_{t}^{(i)}|s_{t-1}^{(i)})}}\left(\nabla\pi_{\theta}(s_{t}^{(i)}|s_{t-1}^{(i)})-\nabla\pi_{\theta_{\text{old}}}(s_{t}^{(i)}|s_{t-1}^{(i)})\right)\cdot\left(r_{i}-\overline{\mu}_{G}\right)}_{\text{\emph{Gradient Error}}}$
$\quad+\underbrace{\frac{1}{|G|}\sum_{i=1}^{|G|}\sum_{t=1}^{T}\bm{1}_{\mathcal{B}(s_{t}^{(i)},\theta,\theta_{\text{old}})}w(s_{t}^{(i)},\theta,\theta_{\text{old}})\nabla\log\pi_{\theta}(s_{t}^{(i)}|s_{t-1}^{(i)})\cdot\left(A_{i}-\frac{r_{i}-\overline{\mu}_{G}}{\overline{\sigma}_{G}}\right)}_{\text{\emph{Sampling Error}}}$
$\quad\underbrace{-\frac{1}{\overline{\sigma}_{G}}\frac{1}{|G|}\sum_{i=1}^{|G|}\sum_{t=1}^{T}\bm{1}_{\mathcal{B}^{c}(s_{t}^{(i)},\theta,\theta_{\text{old}})}w(s_{t}^{(i)},\theta,\theta_{\text{old}})\nabla\log\pi_{\theta}(s_{t}^{(i)}|s_{t-1}^{(i)})\cdot\left(r_{i}-\overline{\mu}_{G}\right)}_{\text{\emph{Clip Error}}}.$ (6)
@@
$\operatorname{\mathbb{E}}_{G}\left[\widetilde{\nabla}J(\theta_{\text{old}})\right]$ $=\operatorname{\mathbb{E}}_{G}\left[\frac{1}{|G|}\sum_{i=1}^{|G|}\left(\nabla\log\left(\prod_{t=1}^{T}\pi_{\theta_{\text{old}}}(s_{t}^{(i)}|s_{t-1}^{(i)})\right)\right)\cdot(r_{i}-\overline{\mu}_{G})\right]$
$=\operatorname{\mathbb{E}}_{s_{T}\sim\pi_{\theta_{\text{old}}}}\left[\left(\nabla\log\operatorname{\mathbb{P}}_{\theta}(s_{T}|s_{0})\right)r(s_{T})\right]-0$
@@
The key insight is that the old policy $\pi_{\theta_{\text{old}}}$ is refreshed every $K$ steps, i.e., $\pi_{\theta_{\text{old}}}\leftarrow\pi_{\theta}$. As a result, the discrepancy between $\pi_{\theta}$ and $\pi_{\theta_{\text{old}}}$ remains small throughout training, allowing the algorithm to perform reliably even with stale gradient estimates. We empirically validate our hypothesis through a controlled ablation experiment. Specifically, we remove the importance sampling mechanism entirely from the DAPO algorithm and, within each inner optimization loop where the old policy $\pi_{\theta_{\mathrm{old}}}$ is held fixed, directly perform updates using the policy gradient estimated at $\pi_{\theta_{\mathrm{old}}}$. This setting isolates the effect of importance sampling and allows us to examine how well GRPO performs when relying solely on stale gradients.

### v2

@@ Abstract
Group Relative Policy Optimization (GRPO), recently proposed by DeepSeek, is a critic free reinforcement learning algorithm for fine tuning large language models. It replaces the value function in Proximal Policy Optimization (PPO) with group normalized rewards, while retaining PPO style token level importance sampling based on an old policy. We show that GRPO update rule in fact estimates the policy gradient at the old policy rather than the current one. However, since the old policy is refreshed every few steps, the discrepancy between the two remains small limiting the impact of this bias in practice. We validate this through an ablation study in which importance sampling is entirely removed, and updates are instead performed using the gradient estimated at a fixed old policy across multiple optimization steps. Remarkably, this simplification results in performance comparable to standard GRPO.
@@ 2.1 Reinforcement Learning Setup
$J(\theta)=\mathbb{E}_{s_{T}\sim\pi_{\theta}}\left[r(s_{T})\right]=\sum_{s_{T}\in\mathcal{S}_{T}}\operatorname{\mathbb{P}}_{\theta}(s_{T}|s_{0})r(s_{T}),$ (2)
@@
$\theta_{n+1}=\theta_{n}+\eta\nabla_{\theta_{n}}\mathcal{J}(\theta_{n}),$
with learning rate sequence $\{\eta\}_{n\geq 1}$. Algorithms like PPO and GRPO build on this principle with various modifications to improve performance.
@@ 2.2 Review of GRPO
$\mathcal{L}_{\text{GRPO}}(\theta,\theta_{\text{old}},\theta_{\text{ref}})$ $=\frac{1}{|G|}\sum_{i=1}^{|G|}\frac{1}{|s_{t}^{(i)}|}\sum_{t=1}^{T}\min\left\{w(s_{t}^{(i)},\theta,\theta_{\text{old}})A_{i},\operatorname{clip}\bigl{(}w(s_{t}^{(i)},\theta,\theta_{\text{old}}),\epsilon_{\text{low}},\epsilon_{\text{high}}\bigr{)}A_{i}\right\}$
$\quad-\beta\mathrm{D}_{\text{KL}}\bigl{(}\pi_{\theta}\|\pi_{\theta_{\text{ref}}}\bigr{)},$ (5)
where $|s_{t}^{(i)}|$ denotes the length of the response $s_{t}^{(i)}$, and the importance ratio $w(s_{t}^{(i)},\theta,\theta_{\text{old}})={\pi_{\theta}(s_{t}^{(i)}\mid s_{t-1}^{(i)})}\Big{/}{\pi_{\theta_{\text{old}}}(s_{t}^{(i)}\mid s_{t-1}^{(i)})},$ and the clipping function
$\operatorname{clip}(x,\epsilon_{\text{low}},\epsilon_{\text{high}})=\begin{cases}1-\epsilon_{\text{low}},&x<1-\epsilon_{\text{low}},\\[4.0pt] x,&1-\epsilon_{\text{low}}\leq x\leq 1+\epsilon_{\text{high}},\\[4.0pt] 1+\epsilon_{\text{high}},&x>1+\epsilon_{\text{high}}.\end{cases}$
In original GRPO, the clipping thresholds used in the surrogate objective are symmetric, i.e., $\epsilon_{\text{low}}=\epsilon_{\text{high}}$. A subsequent study found that using asymmetric clipping ($\epsilon_{\text{low}}\neq\epsilon_{\text{high}}$) can lead to improved empirical performance, and accordingly renamed the modified algorithm as Decouple Clip and Dynamic Sampling Policy Optimization (DAPO) (Yu et al., 2025). Meanwhile, in DAPO, the coefficient ${1}/{|G|}\sum_{i}1/|s_{t}^{(i)}|\sum_{t}(\cdot)$ preceding the loss function is replaced by ${1}/{\sum_{i=1}^{|G|}|s_{T}^{(i)}|}\sum_{i}\sum_{t}(\cdot),$ where $|s_{T}^{(i)}|$ denotes the length of the response $s_{T}^{(i)}$. In fact, since the sequence $\{|s_{T}^{(i)}|\}_{i=1}^{|G|}$ consists of independent and identically distributed random variables, the Law of Large Numbers implies that the summation in the denominator converges to $|G|\cdot\mathbb{E}_{s_{T}\sim\pi_{\theta_{\text{old}}}}[|s_{T}|\mid\mathscr{F}_{\theta_{\mathrm{old}}}]$ as $|G|\to\infty$. Consequently, analyzing this variant only introduces a different sampling error term in the gradient decomposition (shown in Section 3), without posing any essential new theoretical difficulties. Therefore, for the sake of notational simplicity, we adopt the form as original GRPO in this part throughout this paper.
In the remainder of this paper, we do not distinguish between the names DAPO and GRPO.
@@
$\theta_{k+1}=\theta_{k}+\eta\widehat{\nabla}\mathcal{L}_{\text{GRPO}}(\theta_{k},\theta_{\text{old}},\theta_{\text{ref}}),$
where $\eta$ is the learning rate  [^2] Here, $n$ denotes the outer loop index (i.e., the number of major update rounds). Within each inner loop—where the old policy $\pi_{\mathrm{old}}$ remains fixed—we assume the learning rate stays constant. See Algorithm 1 for details on the outer/inner loop structure. and $\widehat{\nabla}\mathcal{L}_{\text{GRPO}}(\theta,\theta_{\text{old}},\theta_{\text{ref}})$ denotes a mini-batch estimate of the gradient $\nabla\mathcal{L}_{\text{GRPO}}(\theta,\theta_{\text{old}},\theta_{\text{ref}})$ which can be written as
$\nabla\mathcal{L}_{\text{GRPO}}(\theta,\theta_{\text{old}},\theta_{\text{ref}})$ $=\underbrace{\frac{1}{|G|}\sum_{i=1}^{|G|}\frac{1}{|s_{T}^{(i)}|}\sum_{t=1}^{T}\nabla\left(\min\left\{w(s_{t}^{(i)},\theta,\theta_{\text{old}})A_{i},\operatorname{clip}\bigl{(}w(s_{t}^{(i)},\theta,\theta_{\text{old}}),\epsilon_{\text{low}},\epsilon_{\text{high}}\bigr{)}A_{i}\right\}\right)}_{\emph{Gradient Term}}$
$\quad\underbrace{-\beta\nabla\left(\mathrm{D}_{\text{KL}}\bigl{(}\pi_{\theta}\|\pi_{\theta_{\text{ref}}}\bigr{)}\right)}_{\emph{Penalty Term}}.$ (6)
@@
Input: reward function $r(\cdot)$; initial parameters $\theta$; reference parameters $\theta_{\mathrm{ref}}$; group size $|G|;$ clipping threshold $\epsilon_{\text{low}},\epsilon_{\text{high}}$; KL coefficient $\beta$;
inner iterations $K$; learning rate $\eta$
@@ 10 $A_{i}\leftarrow\dfrac{r_{i}-\mu_{G}}{\sigma_{G}+\delta}$;
14    $\mathcal{L}\leftarrow\frac{1}{|G|}\sum_{i=1}^{|G|}\frac{1}{|s_{T}^{(i)}|}\sum_{j=1}^{T}\left(\min\left\{w_{i,j}A_{i},\operatorname{clip}\bigl{(}w_{i,j},\epsilon_{\text{low}},\epsilon_{\text{high}}\bigr{)}A_{i}\right\}\;-\beta\text{D}_{\mathrm{KL}}\Bigl{(}\pi_{\theta}\bigl{(}s_{j}^{(i)}\mid s^{(i)}_{j-1}\bigr{)}\|\pi_{\theta_{\mathrm{ref}}}\bigl{(}s_{j}^{(i)}\mid s^{(i)}_{j-1}\bigr{)}\Bigr{)}\right)$;
@@
16    $\theta\leftarrow\theta+\eta\hat{\nabla}\mathcal{L}$;
@@ 3 A Decomposition of GRPO’s Gradient Term
$\quad\text{\emph{Gradient Term}}\mathop{=}^{(*)}\frac{1}{|G|}\sum_{i=1}^{|G|}\frac{1}{|s_{T}^{(i)}|}\sum_{t=1}^{T}\bm{1}_{\mathcal{B}(s_{t}^{(i)},\theta,\theta_{\text{old}})}w(s_{t}^{(i)},\theta,\theta_{\text{old}})\nabla\log\pi_{\theta}(s_{t}^{(i)}|s_{t-1}^{(i)})\cdot A_{i}$
${=}\frac{T_{\theta_{\text{old}}}^{(-1)}}{\overline{\sigma}_{G}}\underbrace{\frac{1}{|G|}\sum_{i=1}^{|G|}\sum_{t=1}^{T}\bm{1}_{\mathcal{B}(s_{t}^{(i)},\theta,\theta_{\text{old}})}\nabla\log\pi_{\theta_{\text{old}}}(s_{t}^{(i)}|s_{t-1}^{(i)})\cdot\left(r_{i}-\overline{\mu}_{G}\right)}_{\widetilde{\nabla}J(\theta_{\text{old}})}$
$\quad+\underbrace{\frac{T_{\theta_{\text{old}}}^{(-1)}}{\overline{\sigma}_{G}}\frac{1}{|G|}\sum_{i=1}^{|G|}\sum_{t=1}^{T}\bm{1}_{\mathcal{B}(s_{t}^{(i)},\theta,\theta_{\text{old}})}\frac{1}{\pi_{\theta_{\text{old}}(s_{t}^{(i)}|s_{t-1}^{(i)})}}\left(\nabla\pi_{\theta}(s_{t}^{(i)}|s_{t-1}^{(i)})-\nabla\pi_{\theta_{\text{old}}}(s_{t}^{(i)}|s_{t-1}^{(i)})\right)\cdot\left(r_{i}-\overline{\mu}_{G}\right)}_{\text{\emph{Gradient Error}}}$
$\quad+\underbrace{\frac{T_{\theta_{\text{old}}}^{(-1)}}{|G|}\sum_{i=1}^{|G|}\sum_{t=1}^{T}\bm{1}_{\mathcal{B}(s_{t}^{(i)},\theta,\theta_{\text{old}})}w(s_{t}^{(i)},\theta,\theta_{\text{old}})\nabla\log\pi_{\theta}(s_{t}^{(i)}|s_{t-1}^{(i)})\cdot\left(A_{i}-\frac{r_{i}-\overline{\mu}_{G}}{\overline{\sigma}_{G}}\right)}_{Sampling\ Error_{1}}$
$+\underbrace{\frac{1}{|G|}\sum_{i=1}^{|G|}\left(T_{\theta_{\text{old}}}^{(-1)}-\frac{1}{|s_{T}^{(i)}|}\right)\sum_{t=1}^{T}\bm{1}_{\mathcal{B}(s_{t}^{(i)},\theta,\theta_{\text{old}})}w(s_{t}^{(i)},\theta,\theta_{\text{old}})\nabla\log\pi_{\theta}(s_{t}^{(i)}|s_{t-1}^{(i)})\cdot A_{i}}_{Sampling\ Error_{2}}$
$\quad\underbrace{-\frac{T_{\theta_{\text{old}}}^{(-1)}}{\overline{\sigma}_{G}}\frac{1}{|G|}\sum_{i=1}^{|G|}\sum_{t=1}^{T}\bm{1}_{\mathcal{B}^{c}(s_{t}^{(i)},\theta,\theta_{\text{old}})}w(s_{t}^{(i)},\theta,\theta_{\text{old}})\nabla\log\pi_{\theta}(s_{t}^{(i)}|s_{t-1}^{(i)})\cdot\left(r_{i}-\overline{\mu}_{G}\right)}_{\text{\emph{Clip Error}}}.$ (7)
In the above expression, we define
$T_{\theta_{\mathrm{old}}}^{(-1)}:=\mathbb{E}_{s_{T}\sim\pi_{\theta_{\text{old}}}}[|s_{T}|^{-1}\mid\mathscr{F}_{\theta_{\mathrm{old}}}],\ \ \overline{\sigma}_{G}:=\mathbb{E}_{s^{(1)},\ s^{(2)},...,\ s^{(|G|)}\sim\pi_{\theta_{\text{old}}}}\left[\sigma_{G}\mid\mathscr{F}_{\theta_{\mathrm{old}}}\right],$
$\overline{\mu}_{G}:=\mathbb{E}_{s^{(1)},\ s^{(2)},...,\ s^{(|G|)}\sim\pi_{\theta_{\text{old}}}}\left[\mu_{G}\mid\mathscr{F}_{\theta_{\mathrm{old}}}\right].$ (8)
@@
$\quad\operatorname{\mathbb{E}}_{s^{(1)},\ s^{(2)},...,\ s^{(|G|)}\sim\pi_{\theta_{\text{old}}}}\left[\widetilde{\nabla}J(\theta_{\text{old}})|\mathscr{F}_{\theta_{\text{old}}}\right]$
$=\operatorname{\mathbb{E}}_{s^{(1)},\ s^{(2)},...,\ s^{(|G|)}\sim\pi_{\theta_{\text{old}}}}\left[\frac{1}{|G|}\sum_{i=1}^{|G|}\left(\nabla\log\left(\prod_{t=1}^{T}\pi_{\theta_{\text{old}}}(s_{t}^{(i)}|s_{t-1}^{(i)})\right)\right)\cdot(r_{i}-\overline{\mu}_{G})\Bigg{|}\mathscr{F}_{\theta_{\text{old}}}\right]$
$=\operatorname{\mathbb{E}}_{s_{T}\sim\pi_{\theta_{\text{old}}}}\left[\left(\nabla\log\operatorname{\mathbb{P}}_{\theta}(s_{T}|s_{0})\right)r(s_{T})|\mathscr{F}_{\theta_{\text{old}}}\right]-0$
@@
The key insight is that the old policy $\pi_{\theta_{\text{old}}}$ is refreshed every $K$ steps, i.e., $\pi_{\theta_{\text{old}}}\leftarrow\pi_{\theta}$. As a result, the discrepancy between $\pi_{\theta}$ and $\pi_{\theta_{\text{old}}}$ remains small throughout training, allowing the algorithm to perform reliably even with stale gradient estimates. We empirically validate our hypothesis through a controlled ablation experiment. Specifically, we remove the importance sampling mechanism entirely from DAPO and, within each inner optimization loop where the old policy $\pi_{\theta_{\mathrm{old}}}$ is held fixed, directly perform updates using the policy gradient estimated at $\pi_{\theta_{\mathrm{old}}}$. This setting isolates the effect of importance sampling and allows us to examine how well GRPO performs when relying solely on stale gradients.